Your banker has analyzed your company account and has suggested that her bank has a cash management package for you. She suggests that with a concentration banking system, your float can be reduced by two days on average. You, of course, are delighted (you’re not sure why), but you do know your average daily collections amount to $370,000. Your opportunity cost of funds is 7 percent. The bank provides this service for $59,000 plus a compensating balance in your current account of $81,500. (A compensating balance is the amount you are required to maintain interest free at that bank.)

a. Is this package worth it?

No Yes

b. By how much? (Negative answer should be indicated by a minus sign.) Annual Saving $
Answer #1

Cost of funds saved due to reducing float by 2 days

=Average daily collection * Float reduced*cost of fund

=370,000*2*7%

=51800

To find the cost of cash management;

cost of cash management =Bank Service fee+cost of funds on compensating balance

=59,000 +(81,500*7%)

=64705

Qa:

No , the package should not be be accepted as the cost saved due to management package is lower than cost of package.

Qb:

Annual loss= 51800 -64705

=-12905 loss
Answer #2

SOLUTION :


Collection in one day = $370000 .

Collection in 2 days = $370000 * 2 = $740000 .

Collection period is reduced by 2 days means cash of $740000 will not be blocked in the system and will be available to the company.


So, 


Saving of interest on this cash-increase in cash availability throughout the year


= 51800 ($).


Cost of getting this system 

= Fees of the bank + loss of interest on compensating balance

= 59000 + 81500*0.07

= 64705 ($)


a.


The package offered by the bank is not worth taking since savings are less than the cost. (ANSWER).


b.


The extent of loss 

= Savings - Cost 

= 51800 - 64705 

= - 12905  ($)  (negative sign indicates loss) (ANSWER)

answered by: Tulsiram Garg